OpenStudy (anonymous):

Rewrite the expression without using a negative exponent 4y^-2

OpenStudy (anonymous):

4/y^2

OpenStudy (unklerhaukus):

x^-1 = 1/x x^-2 = 1/x^2

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

is the question about 4(y)^-2 or (4y)^-2 ?

OpenStudy (anonymous):

Good question. I assumed 4(x^-2) not (4x)^-2